OpenStudy (anonymous):

simplify the expression -3(b-7)

OpenStudy (anonymous):

The expression is already simplified, but its expanded form is 21-3b.

OpenStudy (anonymous):

-3(b-7) =-3b+21 =21-3b

OpenStudy (anonymous):

OK, so how do you get that - exactly???

OpenStudy (anonymous):

-3(b-7) you need to distribute -3b+21

OpenStudy (anonymous):

Thank you so much!